This is driving me insane
The sum of the lengths of any two sides of a triagle must be greater than the third side. If a triagle has one side that is 19 inches and a second side that is 2 inches less than twice the third side, what are the possible lengths for the second and third sides.

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

The sum of the lengths of any two sides of a triagle must be greater than the third side. If a triagle has one side that is 19 inches and a second side that is 2 inches less than twice the third side, what are the possible lengths for the second and third sides.
the question is asking for an answer of :
the second side is between_13 3/4_and 19 inches and the 3rd side is between
_5 1/4_ and _19______.
-----------------
One side = 19 inches
3rd side = x
2nd side = 3x-2
-----------------
EQUATION:
x + 3x-2 >19
4x > 21
x > 5 1/4 in
===================
Cheers,
Stan H.

Answer by checkley71(8403)

X+(2X-2)>19
3X-2>19
3X>21
X>21/3
X>7 for the shorter unknown side.
THEN 2*7-2=14-2=12 or the other unknown side is >12